a game is played with the special fair cubic die which as one red side,two blue sides,and three green sides.the RESULT is the colour f the top side after the die has een rolled.if the is rolled repeatedly,d probability dat the second BLUE RESULT occurs on or b4 d tenth roll,can b expressed in d form (3p-2q)/3rwhere p,q and r are positive integers ,find d value of p2+q2+r2..!!


and d curect ans is 283...!!

Grade:12

2 Answers

Chetan Mandayam Nayakar
13 years ago

probability=10C2/32=5=(32-22)/30, (p,q,r)=(2,2,0) therefore correct answer is 8

'Win exciting gifts by answering the questions on Discussion Forum.'

James Shown
3 years ago
Given data:
1- Red
2-Blue
3-Green sides
Probability of 2 times blue out of 10 rolls = 1 – (Probability of 0 times blue out of 10 + Probability of 1 time blue out of 10)
Probability of 2 times blue out of 10 rolls = 
= 1 – [ ((4/6)^10) + 10c1 (4/6)^9 (2/6)^9]
= (610 – 410 – 410 . 5) / 610
= (610 – 6. 410 )/ 610
= (69 – 410 )/ 69
= (39 – 211 )/ 39
= 3p – 2q / 3r
p=9, q=11, r=9
p2 + q2 + r2 = 283